OpenStudy (anonymous):

Area between curves. Integrate with respect to x or y. Use an approximating rectangle. y=x^1/2, y=1/3x, x=25

OpenStudy (p0sitr0n):

y = 1/(3x) or y =1/3 * x ?

OpenStudy (anonymous):

(1/3)*x

OpenStudy (p0sitr0n):

 guess you need the shaded area \[\int\limits_{0}^{9}\sqrt{x}-\frac{ x }{3}dx\]